OpenStudy (anonymous):

lny=9t-8; solve for y?

OpenStudy (anonymous):

y=e^(9t-8)

ganeshie8 (ganeshie8):

\(\ln y = x => y = e^x\) \(\ln y = 9t-8 => y = ?\)

OpenStudy (anonymous):

i think thats the right answer but how did you get it?

OpenStudy (anonymous):

simple log lindsay \[\log_{a}b=c \] can be also be written as \[a^c=b\]

OpenStudy (anonymous):

thanks! solve 2^x+2^-x=3